OpenStudy (anonymous):

What is the value of r in the equation 4r − 8 = 5r + 12? −20 −1 1 20

OpenStudy (anonymous):

OK get r on 1 side -r-8=12

OpenStudy (anonymous):

Solve for r 8+12=20

OpenStudy (anonymous):

I don't get it giv me a sec

OpenStudy (kendricklamar2014):

\[4r−8=5r+12\] Step 1: Subtract 5r from both sides. \[4r−8−5r=5r+12−5r\] \[−r−8=12\] Step 2: Add 8 to both sides. \[−r−8+8=12+8\] \[−r=20\] Step 3: Divide both sides by -1. \[\frac{ -r }{ -1 } = \frac{ 20 }{ -1 }\] r=? @77777jeannie77777

OpenStudy (anonymous):

-20?

OpenStudy (kendricklamar2014):

Yes that's correct

OpenStudy (anonymous):

Yes it's -20

OpenStudy (anonymous):

yay thanks lol im going to give the medal to @usernamenotfound because @KendrickLamar2014 u are already level 99 :D is that allright?

OpenStudy (anonymous):

i have more questions though :D

OpenStudy (kendricklamar2014):

Yeah thats alright :D

OpenStudy (kendricklamar2014):

I have to go, maybe @usernamenotfound can help you..

OpenStudy (anonymous):

ok :)

OpenStudy (anonymous):

I'm at lunch but I'll help